OpenStudy (anonymous):

The width of the swimming pool is one third of its length. The width of the pool is 15 feet. What is the length of the pool?

OpenStudy (bigpapa1):

isn't this math?

OpenStudy (anonymous):

Either way: Let L be the length of the pool. We know that the width is 1/3 of the length, let W be the width and get W = 1/3 L. We are given that the width is 15, so plug that in for W to get 15 = 1/3 L. Multiply both sides by 3 to get rid of the 1/3 and get 45 = L.

OpenStudy (bigpapa1):

sorry, I'm afraid I can't help you. I wish I could but reading math is like reading another language to me. I don't understand why there are letters in math cause it's supposed to be numbers, you know?

OpenStudy (anonymous):

Yes thank you both so much for your help :)